Find the determinant of the matrix,
A=[1,2,3/-1,2,5/4,0,2]

Answer by Alan3354(69443)

Find the determinant of the matrix,
A=[1,2,3/-1,2,5/4,0,2]
|+1 2 3|
|-1 2 5|
|+4 0 2|
===========
Det = 1*(4-0) - 2*(-2-20) + 3*(0-8)
= 4 + 44 - 24
= 24
---------
Using the minors of the bottom row is easier, it has a zero.
Det = 4*(10-6) + 2*(2+2)
= 16 + 8
= 24